How much water is needed to dilute a mixture of 0.25 moles NaOH and 250 ml H2O so that the solution will have a concentration of 5 ppm?

1
Expert's answer
2021-02-22T05:54:54-0500

5 ppm = 0.000125 M.

250 ml = 0.25 L.

Cm=n/V. = 0.25/0.25 = 1 M.

V= n/Cm= 0.25/0.000125 = 2 000 L.

V(H2O)= V(total) - V(solution) = 2000 - 0.25 = 1999.75 L.